Network status: the start of 2016A
The 2016A semester began on April 1. To those of you who have been granted observing time on the Network, we hope that your observations go well. If you have questions about how to prepare your observing requests, please email our Science Support team.
In the last status report, we mentioned that LCOGT has created a new archive to store its data products. The IPAC archive will remain available until May 31, but we encourage you to retrieve your data from the new archive. To login, use the same username/password that you use for the observatory portal. If you'd like to report a problem with the archive (or make a suggestion), clicking on the "Feedback" link at the bottom of the archive's UI will open a form that you can fill out and submit.
The new semester also marks the deployment of a new data pipeline. We're now processing data through both the new pipeline and the ORAC-DR pipeline. Products from both are available in the archive; products from the new pipeline are indicated with "Reduction Level = Reduced (Experimental)" in the archive interface. Tests of the new pipeline are ongoing, so please alert us if you encounter problems with the reduction procedure. More information on the new pipeline will be made available on the website soon.
All images (raw and pipeline-processed) in the archive are Rice-compressed ("fpacked"). We began compressing all images generated at our observing sites on April 6. The compressed images require less bandwidth to transport off-site and back to LCOGT headquarters for processing.
The Spectral camera on the 2-meter telescope at Siding Spring (FTS) is currently not available for observations. The camera was restored on April 4, but then it warmed-up toward the end of the April 6 night. The FLOYDS (spectrometer) is still available on the 2-meter telescope, but ...
Also at Siding Spring, we've been having intermittent problems with our main site computer. These problems have been driving our IT team crazy, but a consequence for users is that, when the computer crashes, the headers of FITS images can get corrupted. This is a reason why some images made with Siding Spring telescopes aren't yet available. Fixing the site computer is a high-priority task, and we are compiling an inventory of affected images.
At Cerro Tololo, the Sinistro camera (fl03) in Dome B is sometimes producing all-noise images (i.e. only pixels with extremely high and extremely low counts). The problem is currently under investigation.
Share on Twitter
Share on Facebook
Finally, I want to draw your attention to the Insight Astronomy Photographer of the Year 2016
competition. We were asked by the competition coordinator at the Royal Museums Greenwich to make you aware that the competition includes a special prize for images captured with robotic telescopes. The reason for their request: "We strongly feel that they [LCOGT users] are the ideal entrants and would benefit from the opportunity of winning this prize, particularly given the wide press attention the competition receives." For more information, see: http://www.rmg.co.uk/discover/astronomy-photographer-competition
. Entries are due on April 14.